Calculate the area under the curve y=x3 +4x+1 from x=-3 to x=3.

1
Expert's answer
2021-12-06T17:37:50-0500
"y=0=>x^3+4x+1=0"

"x\approx-0.24626617"

"Area=-\displaystyle\int_{-3}^{-0.24626617}(x^3+4x+1)dx"

"+\displaystyle\int_{-0.24626617}^{3}(x^3+4x+1)dx"

"=-[\dfrac{x^4}{4}+2x^2+x]\begin{matrix}
 -0.24626617 \\
 -3
\end{matrix}"

"+[\dfrac{x^4}{4}+2x^2+x]\begin{matrix}
 3\\
 -0.24626617
\end{matrix}"

"=0.1240526+35.25+41.25+0.1240526"

"\approx76.748105 ({units}^2)"